what is...... 24+y=9y

Found 2 solutions by themathtutor2009, mananth:
Answer by themathtutor2009(81)

24+y=9y
24=9y-y
24=8y
24/8=y
3=y
y=3

Answer by mananth(16946)

24+y=9y
Add -y to both side
24+y-y= 9y-y
24=8y
divide both sides by 8
24/8 = 8y/8
3=y
OR y=3